# Low-Latency ML Inference by Grouping Correlated Data Objects and Computation

### v1

@@ Abstract.
ML inference workflows often require low latency and high throughput, yet we lack good options for addressing this need. Techniques that reduce latency in other streaming settings (such as caching and optimization-driven scheduling) are of limited value because ML data dependencies are often very large and can change dramatically depending on the triggering event. In this work, we propose a novel correlation grouping mechanism that makes it easier for developers to express application-specific data access correlations, enabling coordinated management of data objects in server clusters hosting streaming inference tasks. Experiments based on a latency-sensitive ML-based application confirm the limitations of standard techniques while showing that our solution yields dramatically better performance. The proposed mechanism is able to maintain significantly lower and more consistent latency, achieves higher node utilization as workload and scale-out increase, and yet requires only minor changes to the code implementing the application.
conference: Preprint; November 2023; arXiv
@@ 1. Introduction
Latency-sensitive AI-based applications are increasingly common (Chen et al., 2021). For instance, in edge intelligence applications, devices pass captured data through Machine Learning (ML) inference pipelines (Deng et al., 2020). Minimizing latency is the highest priority, although high throughput and efficient platform utilization remain important as secondary goals (Crankshaw et al., 2017). Our work starts with the observation that when these pipelines are deployed onto standard Stream Processing (SP) platforms (Isah et al., 2019), a mismatch occurs. SP systems need low latency but are typically optimized for highly parallel stateless computations and data transformations. ML inference is stateful and often context-sensitive, resulting in data dependencies that can defeat the assumptions made in SP schedulers (Shaowang et al., 2021).
@@
The issue is further complicated by unpredictability. ML models depend on pretrained parameter vectors that can be large. For example, a trajectory-computing task for mobile “agents” in a traffic scene might have one set of models trained for cars, another for bicycles, a third for pedestrians, etc. Among the vehicle models there could be one specialized for taxis, another for emergency responders, and a general-purpose model for all others. The ML logic won’t know which agents are present in the scene (and hence which objects will be needed) until runtime. Caches are too small to hold a copy of every object that might be needed, yet even a single cache miss will leave the task waiting while data is fetched. We are not aware of any SP or ML serving infrastructure with a mix of features matched to these requirements.
To assess the importance of this effect, we ran latency-sensitive ML models on both a local cluster of servers and a public cloud environment using scalable SP server-pool technologies. In all cases we observed frequent pipeline stalls stemming from data accesses that required fetching objects over the datacenter network. These stalls contribute to request backlogs that increase latency, degrade throughput and utilization, and inflate costs.
@@
The correlation grouping mechanism proposed here has two main goals: (i) offer runtime mechanisms whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way; and (ii) leverage this data to significantly reduce latency, particularly as workload increases and the platform scales out. The mechanism consists of three elements: developer-specified logic that will attach a string, called an correlation key, to each new incoming request; developer-specific code to tag stored data objects with correlation keys; and a runtime engine that makes use of correlation keys to optimize object storage and caching decisions. Leveraging correlation groups does require changes to the runtime platform, but involves minimal changes to the ML application. Benefits include enabling proactive collocation of data with computation, coordinated prefetching, and more efficient cache management.
@@
The remainder of this work is organized as follows. Section 2 motivate our proposals by introducing a representative latency-sensitive ML-based application and identifying shortcomings of modern platforms in face of the application data access patterns. This application is used as a running example throughout this paper. We then propose the correlation grouping mechanism in Section 3. Experimental results showing the benefits of the proposed correlation grouping mechanism are reported in Section 4. Next, we support our motivating claims regarding the shortcomings of modern platforms with experimental results reported in Section 5. Related work is presented in Section 6, and we conclude the paper in Section 7.
@@ 2. Motivation
To motivate our effort we first introduce a Real-Time Collision Prediction (RCP) application which is used as a running example in the remainder of the paper. Although we cast it in a somewhat fanciful setting, the workflow embodied in the RCP is representative of a large, important class of latency-sensitive applications (Crankshaw et al., 2020) seen in settings such as factories, robotic warehouses and commercial retail. We then identify limitations of modern SP and ML serving platforms in face of the data access patterns showcased by the RCP application. Finally, we identify trade-offs when addressing these limitations in modern platforms.
2.1. Real-Time Collision Prediction Application
@@
The goal of the application is to continuously predict potential imminent collisions between agents (e.g. cars, pedestrians, cyclists) in traffic intersections with a window of a few seconds. As seen in Figure 1 the RCP pipeline has three steps: multi-object tracking (MOT), trajectory prediction (PRED), and collision detection (CD). The RCP pipeline focuses on the ML, but would be deployed near cameras, signaling danger by flashing large red lights or notifying the hazard warning systems of suitably equipped vehicles. Implementation details are provided in Section 4.1.
Figure 1. Computational graph of the RCP application.
Computational graph of the RCP application.
Figure 1. Computational graph of the RCP application.
@@
The first step, MOT, is responsible for detecting and tracking agents in video frames received from cameras. The computation in this step consists of detecting all agents in each frame and their positions, and then matching each detected agent with its prior instance(s) if any, to determine speed and trajectory. As input, the MOT requires the current frame and the positions and features of agents found in the prior frame. Output is a set of new positions and features for all agents detected in the frame.
@@
A separate instance of the second step, PRED, is triggered for each new position of each agent detected by the MOT step. In this step, the next 12 positions of an agent are predicted based on the past 8 positions of that agent. Thus this step requires the current position of the agent and the previous 7 positions tracked by MOT. The output is thus a trajectory composed of 12 positions.
@@
Each trajectory predicted by PRED triggers a separate instance of the third step, CD. This step matches the received trajectory with all other trajectories predicted so far for the same frame, in order to detect collisions. In addition to the received trajectory, this step also access all available trajectories from the same frame. For each frame, after all instances of CD have been processed, the predicted trajectories of every pair of detected agents will have been evaluated. The output of each instance of this step is a list of collisions.
2.2. Data Access Patterns and Network Overheads
@@
Each step of the RCP application has a different data access pattern. In the MOT step, each request requires a fresh and potentially big data object – features and positions of agents in the previous frame (up to 10MB in our experiments). The nature of the input (a live stream) implies that these objects will only be used once, hence caching would be ineffective. One way to minimize overhead would be to send the object to the node where the next request will be processed. However, the MOT step only learns which objects it requires at runtime, after receiving the request.
@@
PRED requires small but very fresh objects. Although agent positions can be re-used up to 7 times, there are many PRED instances for each frame, thus the overhead from cache misses can add up and become non-negligible in a time-pressured scenario. This overhead grows as more nodes are added to scale out a deployment and cache misses are more common. CD follows a similar rationale: the same trajectory is accessed many times in a short time (the same frame).
We show how significant these overheads can be in the experiments reported in Sections 4 and 5, while more details on modern platforms limitations can be found in Section 6.
2.3. Trade-offs and Limitations
It may seem that we have expressed an impossible objective, but in fact this is not the case. Developers do understand the data dependencies of their ML models, and we will show that they can express them in an simple way. Our central insight is that this is possible, and that the platform can then use the knowledge gained to improve collocation of data and computation. Existing platforms (such as standard cloud frameworks) lack the hooks required to support this capability, in part because of a presumption that such a mechanism would be in tension with other required features such as load balancing and auto-scaling. We will show that with very modest changes to the runtime infrastructure, significant latency improvements are possible even as these dynamic adaptation mechanisms do their job.
To support these claims, we deployed the RCP application twice. We first focused on a platform called Cascade, which is open source and easily extensible. Here the benefits of correlation grouping were dramatic. But then we deployed the identical RCP logic on Microsoft Azure, and ran experiments carefully following recommendations, and using Azure nodes configured for a high quality of service. Azure has fewer extension hooks than Cascade, which prevents us from implementing correlation grouping. Yet following a similar approach yields significant benefit, as shown in Section 5.
3. Correlation Grouping
We now propose the correlation grouping mechanism. As noted in the introduction, the approach has two broad objectives: (i) to offer runtime mechanisms whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way; and (ii) to leverage this data to significantly reduce latency, particularly as workload increases and the platform scales out. In this section we first discuss requirements for the mechanism that arise from the limitations and trade-offs identified in Section 2. Then we introduce the mechanism per se.
@@ 3.1. Requirements
Ideally, developer-oriented mechanisms should be deployment-agnostic: its behavior should be predictable without deployment details. For example, although Apache Spark Resilient Distributed Datasets (RDDs) (Zaharia et al., 2012) offer a way to specify a list of location preferences when storing data, these are treated as hints. The only way to be certain that computation will be collocated with data is to hand-code a manual routing mechanism, using internal APIs to determine which servers have copies of which data items, their current load and request queue length, and resource availability.
Mechanisms such as location preferences in RDD, the partition key in Cosmos DB (Reagan, 2018), and the hash tag in Redis (Redis, 2023) all facilitate collocation of correlated data. For example, data about positions of agents in the same image could be collocated in our RCP application. The cloud runtime environment then introduces further mechanisms intended to collocate  stages of a pipelined computation (for example, inference MOT requests associated with the same RCP video stream) so that a single node will handle them. Other examples of collocating computation include Apache Spark Streaming partitions and Apache Storm stream grouping (Storm, 2023). The problem here is that the data collocation features are decoupled from the computation collocation options, and each service uses its own way of expressing correlations. Developers need a unified mechanism that spans the full deployment and covers both storage placement choices and computational placement decisions.
A mechanism should also be expressive enough to enable the platform to track the data and computation correlations seen at runtime. For example, a single computation may require several data objects, and a single data object may be required by multiple computations. Keeping track of such correlations enables the platform to make optimized decisions regarding data and computation placement. Finally, the mechanism should be lightweight, i.e. the overhead introduced by the mechanism should be negligible and remain constant as the platform scales out.
3.2. Correlation Grouping Mechanism
Our goals and requirements create a complex problem, particularly in modern cloud architectures where one often finds completely independent subsystems specialized for different roles, each with its own design. In particular, storage and computation are generally treated as distinct subsystems, despite the fact that many compute nodes have substantial caches that can be viewed as a component of the storage layer. This observation became the launch point for our correlation grouping mechanism, which groups correlated objects (and hence is a storage-layer feature), yet exposes the locations of these groups so that runtime schedulers can leverage object location data in their decisions.
The next question is how correlation grouping should be expressed. A common approach that we rejected as inflexible assumes that when coding an application, developers can already anticipate the data objects required by each computational stage and declare these dependencies (e.g. using a declarative language or a SQL query) (Zhou et al., 2022a; Makrynioti and Vassalos, 2021; Shaowang et al., 2021). The platform then can parse the specification and create its own internal representation to keep track of which data objects are required by which computations. However, as seen in the RCP application, an ML model might dynamically fetch large objects based on analysis of its inputs. Such model would have an unspecifiable yet latency-critical data dependency.
The alternative approach favored here is to add a label to every request, similar in spirit to hash tags in Redis or partition keys in Azure EHs. The benefit is that labels can be computed dynamically as inputs are classified, so that (for example) an input image showing a taxi could be given a key matching various category-specific ML models and data previously saved in the platform. We call such labels correlation keys: each object has both a unique name (for example, a file pathname or a K/V key) and an correlation key, which would not be unique. Such an approach documents correlations between data and computation: any pair of data/computation with the same correlation key will be understood to be correlated.
The challenge to this approach is that we need to ensure that the mapping between objects and correlation keys can easily be tracked and leveraged across the whole infrastructure. This has to be done in a manner that is lightweight and that can scale to deployments with very large numbers of applications, each with high event rate workflows, and each potentially having a unique correlation key mapping.
@@ 3.2. Correlation Grouping Mechanism
Although it can be implemented in different ways, the proposed correlation grouping mechanism affects two aspects of a platform: the application-level API, and the platform-level runtime engine. In the proposed mechanism, the application-level API provides developers with a way to encode application-specific knowledge, while the platform-level runtime engine makes use of the knowledge provided by developers to improve data/computation collocation, scheduling, prefetching, among other benefits discussed in Section 3.3.
The core of the correlation grouping mechanism is a function $f(r)$, which maps any request $r$ to a set of correlation keys. We abstract different operations in a platform as requests: storing/retrieving a data object is a data request and requesting a computational task to be started is a computation request. Correlation keys keys are labels usually implemented as strings. The application-level API allows developers to provide the correlation grouping function $f$ to the platform. The platform then can apply function $f$ to any request: $f$ extracts information from the request to generate a list of one or more keys. Application-specific knowledge is thus entirely encapsulated in $f$. Note that $f$ will be available throughout the distributed service, and must return the same result for a given request no matter where it is invoked. Moreover, it is often invoked on critical paths, where blocking would be problematic. For example, in Section 4 we discuss an correlation grouping function that uses regular expressions on object keys.
In the RCP application, for example, $f$ can map data requests of past positions of an agent $a$ to a label “agent_a”. Computation requests that start the PRED step to predict the future trajectory of agent $a$ can also be mapped to the same label, letting the platform know at runtime that the corresponding objects and the computation are correlated.
The platform-level runtime engine consists of a set of components distributed across the platform that makes use of the developer-provided function $f$. The platform must guarantee at least that: (i) the location where data objects are stored and/or cached depends on the correlation keys; and (ii) the location where computational tasks are placed depends on the correlation keys. We describe next how correlation grouping satisfies the requirements discussed in Section 3.1.
The proposed mechanism is deployment-agnostic, since function $f$ maps requests to correlation keys according only to application-specific knowledge. The platform is then responsible to decide how to handle requests from each label according to the deployment. Since function $f$ applies to both data and computation requests, the platform-level runtime engine is able to consider both data and computation placement in tandem. Computations and data objects are linked through labels, instead of an explicit static list of input objects or query. Developers have a great deal of flexibility when expressing data access patterns of their application.
@@
Turning to the lightweight requirement, we argue that mechanisms such as hash tags and partition keys are potentially costly: the mapping between object keys and labels must be synchronized across the whole infrastructure, which can put a costly distributed operation such as a database update on the critical path. With correlation grouping, only function $f$ itself must be available on all components of a platform: there is no associated replicated state.

### v2

# Keep Your Friends Close: Leveraging Affinity Groups to Accelerate AI Inference Workflows

@@ Abstract.
AI inference workflows are typically structured as a pipeline or graph of AI programs triggered by events. As events occur, the AIs perform inference or classification tasks under time pressure to respond or take some action. Standard techniques that reduce latency in other streaming settings (such as caching and optimization-driven scheduling) are of limited value because AI data access patterns (models, databases) change depending on the triggering event: a significant departure from traditional streaming. In this work, we propose a novel affinity grouping mechanism that makes it easier for developers to express application-specific data access correlations, enabling coordinated management of data objects in server clusters hosting streaming inference tasks. Our proposals are thus complementary to other approaches such as caching and scheduling. Experiments confirm the limitations of standard techniques, while showing that the proposed mechanism is able to maintain significantly lower latency as workload and scale-out increase, and yet requires only minor code changes.
journalyear: 2025copyright: acmlicensedconference: The 18th ACM International Systems and Storage Conference; September 8–9, 2025; Virtual, Israelbooktitle: The 18th ACM International Systems and Storage Conference (SYSTOR ’25), September 8–9, 2025, Virtual, Israeldoi: 10.1145/3757347.3759129isbn: 979-8-4007-2119-9/2025/09
@@ 1. Introduction
Latency-sensitive AI-based applications are increasingly common (Chen et al., 2021). For instance, in edge intelligence applications, devices pass captured data through AI inference and classification pipelines, frequently combined with continuous learning (Deng et al., 2020). Minimizing latency is the priority, although throughput and resource utilization remain important goals (Crankshaw et al., 2017).
@@
Our work starts with the observation that when these pipelines are deployed onto standard Stream Processing (SP) platforms (Isah et al., 2019), a mismatch occurs. SP systems are typically optimized for highly parallel stateless computations and data transformations. AI inference and classifiers are often stateful and context-sensitive, resulting in data dependencies that can defeat the assumptions made in SP schedulers (Shaowang et al., 2021).
@@
The issue is further complicated by unpredictability. For example, a trajectory-computing task for mobile “actors” in a traffic scene might have different models trained for cars, bicycles, pedestrians, etc. Among the vehicle models there could be one specialized for taxis, another for emergency responders, and a general-purpose model for all others. The AI won’t know which actors are present in the scene (and hence which objects will be needed) until runtime. Caches are too small to hold a copy of every object that might be needed, yet even a single cache miss will leave the task waiting while data is fetched over the network.
Data movement overheads are further amplified when different stages of an AI pipeline are executed in different nodes of a network. Reasons for distributing different stages across multiple nodes include scalability, load balancing, parallelism, and specialized hardware sharing. Scalable storage tends to randomly distribute data objects across the infrastructure. Although schedulers try to minimize data movement when deciding where to execute a pipeline stage, there is no collocation guarantee.
AI developers understand well which collocation dependencies would be beneficial for the latency of their applications. Alas, existing platforms (such as standard cloud frameworks) lack the hooks required for developers to provide feedback. We conjecture that this is partly because of a presumption that such a mechanism would be in tension with core platform features such as load balancing and auto-scaling, and in part because of the concern that the annotation would require significant effort from the developers.
@@
In this work, we propose the affinity grouping mechanism, which has two main goals: (i) to offer standardized, platform-independent annotation method whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way (allowing mechanisms such as auto-scaling and load balancing to keep doing their job); and (ii) leverage this data to significantly reduce latency as workload increases and the platform scales out. The mechanism consists of three elements: developer-specified logic that attaches a string, called an affinity key, to each new incoming request; developer-specific code to tag stored data objects with affinity keys; and a runtime engine that makes use of affinity keys to optimize object and task placement decisions. Leveraging affinity groups does require changes to the runtime platform, but involves minimal changes to the AI application. Benefits include enabling proactive collocation of data with computation, coordinated prefetching, and more efficient cache management. Our proposals are thus complementary to other approaches such as caching and scheduling.
In order to evaluate our proposals, we design and implement a representative latency-sensitive AI-based application consisting of a composition of off-the-shelf AI models. Using this application, two sets of experiments are performed. First, we deploy our representative application on a state-of-the-art stream processing platform targeted at latency-sensitive AI workflows, which is based on a K/V store. Results show that affinity grouping achieves significantly lower and more consistent latency as workload and scale-out increase, compared to the standard object and task placement strategy of the employed platform. All this requires a minimal annotation effort from the developer.
In the second set of experiments, we deploy the same application on a public cloud, Microsoft Azure. The goal is to show evidence of the data movement overheads present in existing SP and AI serving platforms. We observe frequent pipeline stalls stemming from data accesses that required fetching objects over the network. We then modify our application to address the observed overheads, at the cost of a high coupling between application and deployment.
Figure 1. Computational graph of the RCP application.
[…]
Computational graph of the RCP application.
This work makes the following contributions:
  • We identify inefficiencies of modern platforms to collocate data when deploying a representative latency-sensitive AI-based application that comprises off-the-shelf AI models. We support this claim by investigating the application performance on a public cloud.
  • To address the identified issue, we propose affinity grouping, an easy-to-use mechanism that requires no developer knowledge of deployment/environment details, and yet enables platforms to achieve effective collocation of data and compute.
  • An implementation of the affinity grouping mechanism on a state-of-the-art AI serving platform based on a K/V store is described and evaluated.
@@
The remainder of this work is organized as follows. Section 2 motivate our proposals by introducing a representative latency-sensitive AI-based application and identifying shortcomings of modern platforms in face of the application data access patterns. This application is used as a running example throughout this paper. We then propose the affinity grouping mechanism in Section 3. Experimental results showing the benefits of the proposed affinity grouping mechanism on a local cluster are reported in Section 4. A series of experiments on Microsoft Azure that support our motivating claims are reported in Section 5. Related work is presented in Section 6, while Section 7 discuss the generality of our approach. We conclude the paper in Section 8.
@@ 2. Motivation
We first introduce a Real-Time Collision Prediction (RCP) application that is used as a running example in this paper. RCP is representative of a large, important class of latency-sensitive applications (Crankshaw et al., 2020) seen in settings such as factories, robotic warehouses and commercial retail. We then identify limitations of modern SP and AI serving platforms in face of the data access patterns showcased by the RCP application.
2.1. Real-Time Collision Prediction
@@
The RCP application continuously sense imminent collisions between actors (e.g. cars, pedestrians) in traffic intersections with a window of a few seconds. It consists of a composition of off-the-shelf AI models (trained for different tasks), resulting in a pipeline that runs in an edge datacenter, near traffic cameras. We focus on the AI inference pipeline, but examples of reactions to the pipeline output could include flashing red lights or notifying suitably equipped vehicles. As seen in Figure 1 the RCP pipeline has three steps: multi-object tracking (MOT), trajectory prediction (PRED), and collision detection (CD). Implementation details are provided in Section 4.1.
@@
The first step, MOT, is responsible for detecting and tracking actors in video frames received from cameras. The computation in this step consists of detecting all actors in each frame and their positions, and then matching each detected actor with its prior instance(s) if any, to determine speed and trajectory. As input, the MOT requires the current frame and the positions and features of actors found in the prior frame. Output is a set of new positions and features for all actors detected in the frame.
@@
A separate instance of the second step, PRED, is triggered for each new position of each actor detected by MOT. PRED predicts the next $q$ positions of an actor based on the past $p$ positions of that actor. Thus this step requires the current position of the actor and the previous p1p-1 positions tracked by MOT. The output is thus a trajectory composed of $q$ positions. In our implementation $p=8$ and $q=12$, since the pretrained model employed was trained with those values.
@@
Each trajectory predicted by PRED triggers a separate instance of the third step, CD. This step matches the received trajectory with all other trajectories predicted so far for the same frame, in order to detect collisions. In addition to the received trajectory, this step also access all available trajectories from the same frame. For each frame, after all instances of CD have been processed, the predicted trajectories of every pair of detected actors will have been evaluated. The output of each instance of this step is a list of collisions.
2.2. RCP Distributed Deployment
Earlier, we noted that applications such as our RCP example lend themselves to a distributed deployment. What we failed to say is that on many platforms, doing so would be a big departure from what the platform supports. In fact, today’s most common approach for implementing an application like the RCP is to combine the whole pipeline into a monolithic application, then run it on one multi-core server, selected by a load balancer. In order to scale and keep up in real-time, load balancing tends to randomly spray requests over nodes. Although this approach avoids overheads associated with transitioning from one step to the next (since all steps happen in the same process), with complex tasks such as the RCP, it could easily exceed what one server can handle.
Consecutive invocations of the pipeline for frames from the same camera have a data dependency. In addition, multiple steps (MOT and PRED) require GPU accelerators. As a consequence, a frame $k$ must wait until the whole pipeline for frame k1k-1 finishes processing and make the necessary input data available. Even frames from other cameras will need GPU resources to be freed. Furthermore, multiple instances of PRED, one for each actor detected by MOT in the same frame, can be executed in parallel. However, in the monolithic approach they would bottleneck, significantly increasing end-to-end latency for that frame. Finally, certain steps may require more resources than others. In the monolithic approach, scaling individual steps becomes a challenge. For example, in our experiments the PRED step presented the heaviest workload, since each frame had up to 49 actors, and thus this step required more resources.
Thus a job like RCP lends itself to multi-server parallelism: now the scheduler can place different steps on different servers, taking care not to overload any single server. However, latency will still be sensitive to data movement overheads related to the data access patterns of each step. We further discuss such overheads next.
2.3. RCP Data Access Patterns and Overheads
When we deploy an AI pipeline on multiple servers, load balancing will often run each new task on a server picked at random among lightly loaded machines. Yet because AI models and data dependencies can be very large, the new task then pauses to fetch data over the network. In real-time applications such as the RCP, standard solutions such as caching are of limited value due to data freshness. We discuss next the extent of this effect, and how developer-provided knowledge can help the platform.
@@ 2.3. RCP Data Access Patterns and Overheads
In the MOT step, each request requires a fresh and potentially big data object – features and positions of actors in the previous frame (up to 10MB in our experiments). The nature of the input (a live stream) implies that these objects will only be used once, hence caching would be ineffective, since cached objects will never be reused. One way to minimize this overhead would be to send the object to the node where the next request will be processed. However, the MOT step only learns which objects it requires at runtime, after receiving the request.
@@
PRED requires small but very fresh objects. Actor positions can be re-used only up to 7 times, and there are many PRED instances for each frame (one for each actor detected). Thus the overhead from cache misses can add up and become non-negligible in a time-pressured scenario, as such instances are placed on different servers by the load balancer. This overhead grows as more nodes are added to scale out and cache misses are more common. Furthermore, which objects will be necessary are only known at runtime, after MOT is finished. CD follows a similar rationale: the same trajectory is accessed many times in a short time (the same frame), and the actual workload is not predictable.
Despite the freshness and the unpredictability of data access, the developer of the RCP application knows the mapping between future requests and objects stored previously. A platform can benefit from this knowledge and collocate correlated objects and requests proactively, slashing data movement overheads. We achieve that with the affinity grouping mechanism, proposed next. Later, we show how significant the overheads discussed above can be in the experiments reported in Sections 4 and 5.
3. Affinity Grouping
In this section, we start with an overview of how data and tasks are processed in modern SP and AI serving platforms. We then discuss requirements for the mechanism arising from the challenges identified in Section 2. Finally, we introduce and discuss the affinity grouping mechanism.
3.1. Execution and Data Flow Model
We now describe a general execution and data flow model for a typical platform that hosts applications such as the RCP. Our proposals are related to the location at which data is stored/retrieved, and computational tasks are executed. A location refers to a network endpoint (e.g. a server node) that holds resources such as CPUs, GPUs, hard disks, and memory. A platform consists of multiple of such resources distributed across an edge/cloud datacenter network.
Platforms can be divided in two main subsystems that constitute the platform-level runtime engine: storage and compute. As an application runs, data objects are stored and retrieved (e.g. actors positions in RCP), and computational tasks are initiated (e.g. a step of the RCP pipeline). In both cases, the corresponding subsystem needs to make placement decisions. The storage subsystem contains a data scheduling component responsible for deciding on the location a data object is to be stored at or retrieved from. Similarly, the compute subsystem has a task scheduling component that decides upon the location at which a computational task is started. Data objects and/or computational tasks are said to be collocated when the corresponding scheduling components place them at the same location.
In addition, a platform also provides developers with an application-level API, used to interact with the platform-level runtime engine. Garbage collection, i.e. how data objects are cleaned from storage when they are not needed (e.g. actor positions in RCP are not needed any longer after some seconds), is orthogonal to this work and thus out of scope.
@@ 3.2. Requirements
Introduction of feedback by the application developer affects two aspects of a platform: the application-level API, and the platform-level runtime engine. The API provides developers with a way to encode application-specific knowledge about correlations, while the platform-level runtime engine makes use of the knowledge provided by developers to improve data/computation collocation, scheduling, prefetching, among other benefits discussed in Section 3.4.
Ideally, the method of specifying correlations through the application-level API exposed by the platform should be deployment-agnostic: the developer logic should not be intertangled with specific deployment details. For example, the only way to be certain that a computation in Apache Spark Resilient Distributed Datasets (RDDs) (Zaharia et al., 2012) will be collocated with data is to manually code a routing mechanism, using internal RDDs APIs to determine which servers have copies of which data items. A similarly complicated proprietary approach is required to ensure collocation in Microsoft Azure, as we show in Section 5. While RDDs offers a way to specify a list of location preferences when storing data, these are only treated as hints, thus not providing a guarantee.
The method of specifying correlations should also be expressive enough to allow the platform to capture the data and computation correlations at runtime. For example, a single computation may require several data objects, and a single data object may be required by multiple computations. A common approach that we reject as inflexible assumes that when coding an application, developers can already anticipate the data objects required by each computational stage and statically declare these dependencies (e.g. using a declarative language or SQL) (Zhou et al., 2022a; Makrynioti and Vassalos, 2021; Shaowang et al., 2021). The platform then can parse the specification and create its own internal representation to keep track of which data objects are required by which computations. However, as seen in the RCP application, an AI model might dynamically fetch large objects based on analysis of its inputs. Such model would have an unspecifiable yet latency-critical data dependency.
Mechanisms such as location preferences in RDD (Zaharia et al., 2012), the partition key in Cosmos DB (Reagan, 2018), and the hash tag in Redis (Redis, 2023) all facilitate collocation of correlated data. For example, positions of actors in the same image could be collocated in our RCP application. The cloud runtime environment then introduces further mechanisms intended to collocate stages of a pipelined computation (for example, MOT inference requests associated with the same RCP video stream) so that a single node handles them. Other examples of collocating computation include Apache Spark Streaming partitions and Apache Storm stream grouping (Storm, 2023). The problem here is that the data collocation features are decoupled from the computation collocation options. Developers need a unified mechanism that spans the full deployment and covers both storage placement choices and computational placement decisions. Furthermore, it is important that the specification method is easy-to-use for application developers, i.e. only minimal changes to the application code should be necessary. Modern APIs for developing AI-based applications generally satisfy this requirement.
@@
Finally, the mechanism should be lightweight: the engine should support specified collocations efficiently at runtime, i.e. the overhead introduced by the implementation within the engine should be negligible and remain constant as the platform scales out. Mechanisms such as Redis hash tags and Azure partition keys are potentially costly: the mapping between object keys and labels must be synchronized across the whole infrastructure, which can put a costly distributed operation such as a database update on the critical path.
Each of the requirements described above is not particularly challenging by itself. The challenge comes when trying to satisfy all requirements together. For instance, the requirement of expressiveness may be in conflict with ease-of-use: providing more information may also increase the complexity of the API for developers. Similarly, designing a unified mechanism may also increase the complexity for developers, and also for the platform to keep track and synchronize correlation information across the infrastructure. Moreover, a deployment-agnostic mechanism may not be expressive enough, since deployment details have to be abstracted away.
3.3. Affinity Grouping Mechanism
Our goals and requirements create a complex problem, particularly in modern cloud architectures where one often finds completely independent subsystems specialized for different roles, each with its own design. In particular, storage and computation are generally treated as distinct subsystems, despite the fact that many compute nodes have substantial caches that can be viewed as a component of the storage layer. This observation became the launch point for our affinity grouping mechanism, which groups correlated objects (and hence is a storage-layer feature), yet exposes the locations of these groups so that runtime schedulers can leverage object location data in their decisions.
In affinity grouping, each data object and computational task has a label, similar in spirit to hash tags in Redis or partition keys in Azure EHs. The critical difference is that labels can be computed dynamically at runtime as inputs are classified. For example, an input image showing a taxi could be given a label matching various category-specific AI models and data previously saved in the platform. We call such labels affinity keys: each object and task has both a unique name (for example, a file pathname or a K/V key) and an affinity key, which would not be unique. Such an approach allows the platform to derive correlations between data and computation: any data objects and a task with the same affinity key will be regarded as requiring correlation.
The core of the proposed mechanism is a function $f(d)$, which maps a descriptor $d$ to an affinity key. A descriptor contains metadata about a data object (to be stored or retrieved) or a computational task (to be initiated). Affinity keys are labels that can be implemented e.g. as strings. The application-level API allows developers to provide the affinity grouping function $f$ to the platform. When a placement decision for a data object or a computational task needs to be made, the platform then can apply function $f$: the function extracts information from the given descriptor to generate an affinity key. Application-specific knowledge is thus entirely encapsulated in $f$. Note that $f$ will be available throughout the distributed service, and must return the same result for a given descriptor no matter where it is invoked. Moreover, it is often invoked on critical paths, where blocking would be problematic. For example, in Section 4 we discuss an affinity grouping function that uses regular expressions.
The platform-level runtime engine must guarantee that the location where data objects are stored and/or cached, and computational tasks are placed, depends on affinity keys. In the RCP application, for example, $f$ can map data objects of past positions of an actor $a$ to a label “actor_a”. The computation task, corresponding to the PRED step, predicting the future trajectory of actor $a$ can also be mapped to the same label, letting the platform know at runtime that the corresponding objects and the computation are correlated.
Regarding the requirements defined above, the proposed mechanism is deployment-agnostic, since function $f$ maps descriptors to affinity keys according only to application-specific knowledge. The platform is then responsible to decide how to handle objects and tasks from each label according to the deployment. The mechanism is unified, since function $f$ applies to placement decisions on both storage and compute subsystems. Tasks and objects are linked through labels, instead of an explicit static list of input objects or query. Developers have a great deal of flexibility when expressing data access patterns of their application (i.e. expressive). Turning to the efficiency requirement, only function $f$ itself must be available on all components of a platform: there is no associated replicated state. Compliance with the ease-of-use requirement, however, depends on the specific implementation. Section 4 describes an implementation in which developers only need to provide a regular expression.